At present, john's age is 30% of his father's age. Twenty years from now, john's age will be 58% of his father's age. How old are they?
Answer by macston(5194)

.
J=John's current age; F=father's current age
.
J=0.30F Use this to substitute for J.
J+20=0.58(F+20) Substitute for J.
0.30F+20=0.58F+11.6
8.4=0.28F
30=F ANSWER 1:Father is 30 years old.
.
J=0.30F=(0.30)30=9 ANSWER 2: John is 9 years old.
.
CHECK:
J+20=0.58(F+20)
9+20=0.58(30+20)
29=0.58(50)
29=29